# Next-order asymptotics for the volume of Schatten balls

### v1

@@ 2. Proofs
In the following, we provide the proofs of Theorems 1 and 2. First, in Lemma 7 we prove Hölder regularity of the Ullman density. Then Theorem 2 is a direct consequence of LS17 . Subsequently, in Lemma 8 we recall a Weyl type integration formula providing the link between $\mathrm{vol}(\mathbb{B}_{p,\beta}^{n})$ and $Z_{n,p,\beta}$. Finally, we deduce Theorem 1 from Theorem 2 with the help of a known asymptotic expansion in Lemma 10.
Corollary 1.1 in LS17 states that the asymptotics in Theorem 2 hold, i.e.
@@
It is known from [Thm. 1.34]DKM98 that if $V\in C^{3}$, that is if $p\geq 3$, then the density of the equilibrium measure is Hölder continuous of order $\frac{1}{2}$ and behaves like $\sqrt{1-|x|}$ near the boundary of its support. To include smaller values of $p$, we verify the following statement directly.
@@
Let $p\geq\frac{3}{2}$. Then $f_{p}$ is Hölder continuous of order $\frac{1}{2}$ on [1,1][-1,1] and there exist positive constants $c$ and $C$ such that
@@
Thus, in this case we obtain that $f_{p}$ is Hölder continuous of order p1p-1 at 0. For $p\geq 2$, a similar argument shows that $f_{p}$ is continuously differentiable at 0 with $f_{p}^{\prime}(0)=0$. In conclusion, if $p\geq\frac{3}{2}$, the function $f_{p}$ is Hölder continuous of order $\frac{1}{2}$ on [1,1][-1,1]. This completes the proof. ∎
@@
Let $p\geq\frac{3}{2}$. In order to apply (LS17, , Cor. 1.1) to $V(x)=v_{p}|x|^{p}$, we verify assumptions (H1)-(H5) there. The function $V$ is lower semi-continuous, bounded from below, everywhere finite and grows polynomially as $|x|\to\infty$. Thus, assumptions (H1) and (H2) in LS17 are satisfied. By Lemma 7 the density $f_{p}$ of the equilibrium measure $\mu_{p}$ is Hölder continuous of order $\frac{1}{2}$ on its support $\Sigma=[-1,1]$. Thus, assumption (H3) is satisfied. For assumption (H4) we note that $\Sigma$ is connected and its boundary is given by $\{-1,1\}$. By Lemma 7, the density $f_{p}$ behaves like $c\sqrt{1-|x|}$ in neighborhoods of the boundary points. Thus, assumption (H4) is satisfied. Assumption (H5) is satisfied since $x\mapsto\exp(-c(|x|^{p}-\ln|x|))$ is integrable on $\mathbb{R}$ for every $c>0$. Thus, we can apply (LS17, , Cor. 1.1) to obtain (14). Finally, the constant $C(\beta)$ can be obtained from the known full expansion of the partition function in the Gaussian case $p=2$, see e.g. (BG24, , Sec. 7). ∎
For the proof of Theorem 1 we provide the following well-known link between Schatten $p$-balls and $\beta$-ensembles. It is a combination of a Weyl type integration formula, see AGZ10 ; SR84 , with a classical trick in convexity going back to Dir39 for computing the volume of the unit ball of $\ell_{p}^{n}$, see DFG+23 ; GP07 ; KMP98 . Recall that $d_{n}=\frac{\beta}{2}n^{2}+(1-\frac{\beta}{2})n$ and $\mathbb{B}_{p}^{n}=\{x\in\mathbb{R}^{n}\colon\sum_{i=1}^{n}|x_{i}|^{p}\leq 1\}$.
@@
$\mathrm{vol}(\mathbb{B}_{p,\beta}^{n})$ $=c_{n}\int_{\mathbb{B}_{p}^{n}}\prod_{1\leq i<j\leq n}|x_{i}-x_{j}|^{\beta}\mathrm{d}x_{1}\cdots\mathrm{d}x_{n}$
@@
$\mathrm{vol}(\mathbb{B}_{\infty,\beta}^{n})=c_{n}\int_{[-1,1]^{n}}\prod_{1\leq i<j\leq n}|x_{i}-x_{j}|^{\beta}\mathrm{d}x_{1}\cdots\mathrm{d}x_{n}.$
@@
where $a_{\beta}\in\mathbb{R}$ is some constant.
@@
where $\Gamma_{2}(x;\frac{2}{\beta},1)$ is a particular instance of the Barnes double Gamma function, see BG24 ; Spr09 . Adapting (BG24, , eq. (7.17)), the following asymptotic expansion holds
@@
where $a_{\beta}^{\prime}\in\mathbb{R}$ is some constant. Combined with Stirling’s approximation $\ln n!=n\ln n-n+\frac{1}{2}\ln(2\pi n)+o(1)$ we deduce the claimed expansion. ∎
@@
Let $\beta\in\{1,2,4\}$ and $p\geq\frac{3}{2}$. Recall that $d_{n}=\frac{\beta}{2}n^{2}+(1-\frac{\beta}{2})n$. By Lemma 8 it holds that
@@
$\ln\Gamma\Big(1+\frac{d_{n}}{p}\Big)$ $=\frac{d_{n}}{p}\Big(\ln\frac{d_{n}}{p}-1\Big)+o(n)$
$=\frac{d_{n}}{p}\Big(2\ln n+\ln\frac{\beta}{2p}-1+\Big(1-\frac{\beta}{2}\Big)\frac{2}{\beta n}\Big)+o(n).$ (19)
@@
$\frac{d_{n}}{p}\ln\Big(\frac{\beta nv_{p}}{2}\Big)-\ln\Gamma\Big(1+\frac{d_{n}}{p}\Big)=$ $-\frac{d_{n}}{p}\Big(\ln n-\ln(pv_{p})-1+\Big(1-\frac{\beta}{2}\Big)\frac{2}{\beta n}\Big)$
@@ Acknowledgement
Funded by the Deutsche Forschungsgemeinschaft (DFG, German Research Foundation) through the SPP 2265 Random Geometric Systems and under Germany’s Excellence Strategy EXC 2044/2 - 390685587, Mathematics Münster: Dynamics-Geometry-Structure. This research was funded in whole or in part by the Austrian Science Fund (FWF) [Grant DOI: 10.55776/J4777]. For open access purposes, the author has applied a CC BY public copyright license to any author-accepted manuscript version arising from this submission.

### v2

@@ 2. Proofs
In the following, we provide the proofs of Theorems 1 and 2. First, in Lemma 6 we prove Hölder regularity of the Ullman density. Then Theorem 2 is a direct consequence of Leblé and Serfaty (2017). Subsequently, in Lemma 7 we recall a Weyl type integration formula providing the link between $\mathrm{vol}(\mathbb{B}_{p,\beta}^{n})$ and $Z_{n,p,\beta}$. Finally, we deduce Theorem 1 from Theorem 2 with the help of a known asymptotic expansion in Lemma 9.
Corollary 1.1 in Leblé and Serfaty (2017) states that the asymptotics in Theorem 2 hold, i.e.
@@
It is known from (Deift et al., 1998, Thm. 1.34) that if $V\in C^{3}$, that is if $p\geq 3$, then the density of the equilibrium measure is Hölder continuous of order $\frac{1}{2}$ and behaves like $\sqrt{1-|x|}$ near the boundary of its support. To include smaller values of $p$, we verify the following statement directly.
@@
Let $p>1$. Then $f_{p}$ is Hölder continuous of order $\min\{p-1,\frac{1}{2}\}$ on [1,1][-1,1]. Moreover, it is Hölder continuous of order $\frac{1}{2}$ in neighborhoods of 1-1 and $1$ and there exist positive constants $c$ and $C$ such that
@@
Thus, in this case we obtain that $f_{p}$ is Hölder continuous of order p1p-1 at 0. For $p\geq 2$, a similar argument shows that $f_{p}$ is continuously differentiable at 0 with $f_{p}^{\prime}(0)=0$. In conclusion, if $p>1$, the function $f_{p}$ is Hölder continuous of order $\min\{p-1,\frac{1}{2}\}$ on [1,1][-1,1]. This completes the proof. ∎
@@
Let $p>1$. In order to apply (Leblé and Serfaty, 2017, Cor. 1.1) to $V(x)=v_{p}|x|^{p}$, we verify assumptions (H1)-(H5) there. The function $V$ is lower semi-continuous, bounded from below, everywhere finite and grows polynomially as $|x|\to\infty$. Thus, assumptions (H1) and (H2) in Leblé and Serfaty (2017) are satisfied. By Lemma 6 the density $f_{p}$ of the equilibrium measure $\mu_{p}$ is Hölder continuous of order $\min\{p-1,\frac{1}{2}\}$ on its support $\Sigma=[-1,1]$ (and Hölder continuous of order $\frac{1}{2}$ away from zero). Thus, assumption (H3) is satisfied. For assumption (H4) we note that $\Sigma$ is connected and its boundary is given by $\{-1,1\}$. Further, by Lemma 6, the density $f_{p}$ behaves like $c\sqrt{1-|x|}$ in neighborhoods of the boundary points and is Hölder continuous of order $\frac{1}{2}$ there. Thus, assumption (H4) is satisfied. Assumption (H5) is satisfied since $x\mapsto\exp(-c(|x|^{p}-\ln|x|))$ is integrable on $\mathbb{R}$ for every $c>0$. Thus, we can apply (Leblé and Serfaty, 2017, Cor. 1.1) to obtain (15). Finally, the constant $C(\beta)$ can be obtained from the known full expansion of the partition function in the Gaussian case $p=2$, see e.g. (Borot and Guionnet, 2024, Sec. 7). ∎
For the proof of Theorem 1 we provide the following well-known link between Schatten $p$-balls and $\beta$-ensembles. It is a combination of a Weyl type integration formula, see Anderson et al. (2010); Saint Raymond (1984), with a classical trick in convexity going back to Dirichlet (1839) for computing the volume of the unit ball of $\ell_{p}^{n}$, see Dadoun et al. (2023); Guédon and Paouris (2007); König et al. (1998). Recall that $d_{n}=\frac{\beta}{2}n^{2}+(1-\frac{\beta}{2})n$ and $\mathbb{B}_{p}^{n}=\{x\in\mathbb{R}^{n}\colon\sum_{i=1}^{n}|x_{i}|^{p}\leq 1\}$.
@@
$\mathrm{vol}(\mathbb{B}_{p,\beta}^{n})$ $=c_{n}\int_{\mathbb{B}_{p}^{n}}\prod_{1\leq i<j\leq n}|x_{i}-x_{j}|^{\beta}\mathrm{d}x$
@@
$\mathrm{vol}(\mathbb{B}_{\infty,\beta}^{n})=c_{n}\int_{[-1,1]^{n}}\prod_{1\leq i<j\leq n}|x_{i}-x_{j}|^{\beta}\mathrm{d}x.$
@@
where $a_{\beta}\in\mathbb{R}$ is some constant with $a_{2}=\frac{1}{2}\ln 2\pi-\zeta^{\prime}(-1)$ with $\zeta$ being the Riemann zeta function.
@@
where $\Gamma_{2}(x;\frac{2}{\beta},1)$ is a particular instance of the Barnes double Gamma function, see Borot and Guionnet (2024); Spreafico (2009). Adapting (Borot and Guionnet, 2024, eq. (7.17)), the following asymptotic expansion holds
@@
where $a_{\beta}^{\prime}\in\mathbb{R}$ is some constant with $a_{2}^{\prime}=\frac{1}{2}\ln 2\pi-\zeta^{\prime}(-1)$. Combined with Stirling’s approximation $\ln n!=n\ln n-n+\frac{1}{2}\ln(2\pi n)+o(1)$ we deduce the claimed expansion. ∎
@@
Let $\beta\in\{1,2,4\}$ and $p>1$. Recall that $d_{n}=\frac{\beta}{2}n^{2}+(1-\frac{\beta}{2})n$. By Lemma 7 it holds that
@@
$\ln\Gamma\Big(1+\frac{d_{n}}{p}\Big)$ $=\frac{d_{n}}{p}\Big(\ln\frac{d_{n}}{p}-1\Big)+\frac{1}{2}\ln\Big(\frac{2\pi d_{n}}{p}\Big)+o(1)$
$=\frac{d_{n}}{p}\Big(2\ln n+\ln\frac{\beta}{2p}-1+\Big(1-\frac{\beta}{2}\Big)\frac{2}{\beta n}\Big)+\frac{1}{2}\ln\Big(\frac{\pi\beta n^{2}}{p}\Big)+o(1).$ (20)
@@
$\frac{d_{n}}{p}\ln\Big(\frac{\beta nv_{p}}{2}\Big)-\ln\Gamma\Big(1+\frac{d_{n}}{p}\Big)=$ $-\frac{d_{n}}{p}\Big(\ln n-\ln(pv_{p})-1+\Big(1-\frac{\beta}{2}\Big)\frac{2}{\beta n}\Big)+o(n)$
@@ Acknowledgement
The author is grateful to Thomas Leblé for a helpful discussion regarding the case $p=1$. Funded by the Deutsche Forschungsgemeinschaft (DFG, German Research Foundation) through the SPP 2265 Random Geometric Systems and under Germany’s Excellence Strategy EXC 2044/2 - 390685587, Mathematics Münster: Dynamics-Geometry-Structure. This research was funded in whole or in part by the Austrian Science Fund (FWF) [Grant DOI: 10.55776/J4777]. For open access purposes, the author has applied a CC BY public copyright license to any author-accepted manuscript version arising from this submission.